OpenStudy (anonymous):

perform the indicated operation: z^2-y^2/a^2-b^2 ÷ z+y/a-b

OpenStudy (anonymous):

\[\frac{z^2-y^2}{a^2-b^2} ÷ \frac{z+y}{a-b}\] \[=\frac{(z-y)(z+y)}{(a-b)(a+b)} \div \frac{(z+y)}{(a-b)}\] \[=\frac{(z-y)(z+y)}{(a-b)(a+b)} \times\frac{(a-b)}{(z+y)}\] \[=\frac{(z-y)}{(a+b)}\]